OpenStudy (anonymous):

Given the function f(x) = 4(x+3) − 5, solve for the inverse function when x = 3.

OpenStudy (anonymous):

First of all, f(x) is basically your y coordinate for a particular x coordinate or value of x. So you can write y=4(x+3)-5.

OpenStudy (gorv):

f=4(x-3)-5 f+5=4*(x-3) (f+5)/4=x-3 (f+5)/4 +3 = x now interchange x anh f (x+5)/4 +3 = f this is ur inverse function

OpenStudy (gorv):

inverse=(x+5)/4 +3

OpenStudy (anonymous):

Just as gorv did, you basically interchange x and y or write x in terms of y. That, will be your inverse function. Remember inverse always interchanges the x and y. Now, you just plug in the value of x to get the required answer.

OpenStudy (gorv):

@silversurfer ...your approch is also right i forgot to look at that ...as i ws in hurry lolll

OpenStudy (anonymous):

do you know the answer?